OpenStudy (anonymous):

Diane has $1.60 in dimes and nickels. She has a total of 20 coins. How many of each kind does she have? @iambatman @kropot72

OpenStudy (anonymous):

x+y=20 .10x+.05 or something like that lol

OpenStudy (anonymous):

x+y = 20 5x+10y=160

OpenStudy (anonymous):

Well close, but you're looking for how many of each kind does she have.

OpenStudy (anonymous):

ha! Nailed it!! lol

OpenStudy (anonymous):

ok, 12 & 8!

OpenStudy (anonymous):

I didn't do it, but I believe in you haha :D and yes the 12 should be nickles and 8 should be dimes.

OpenStudy (anonymous):

lol ok, i'm done with math, ugh...didn't finish, but still got some points :P